### v2

# Nonasymptotic Convergence Rates for Plug-and-Play Methods With MMSE Denoisers

Abstract
It is known that the minimum-mean-squared-error (MMSE) denoiser under Gaussian noise can be written as a proximal operator, which suffices for asymptotic convergence of plug-and-play (PnP) methods but does not reveal the structure of the induced regularizer or give convergence rates. We show that the MMSE denoiser corresponds to a regularizer that can be written explicitly as an upper Moreau envelope of the negative log-marginal density, which in turn implies that the regularizer is $1$-weakly convex. Using this property, we derive (to the best of our knowledge) the first sublinear convergence guarantee for PnP proximal gradient descent with an MMSE denoiser. We validate the theory with a one-dimensional synthetic study that recovers the implicit regularizer. We also validate the theory with imaging experiments (deblurring and computed tomography), which exhibit the predicted sublinear behavior.
I Background
In this work, we focus on the setting of linear inverse problems, where $x\in\mathbb{R}^{n}$ is the true signal, $y\in\mathbb{R}^{m}$ is the observed signal obtained by
[…]
for linear forward operator $A\in\mathbb{R}^{m\times n}$ and measurement noise $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$. The task of recovering $x$ from $y$ is extensively studied in statistics and signal processing. A special case is the denoising problem, obtained when $A=I$:
[…]
Although simpler, denoising is directly useful in solving general inverse problems, often as an intermediate step.
Recovering $x$ from $y$ is often posed as the minimization of a regularized objective function
[…]
where $f$ is a data-fidelity term (typically $f(x)\coloneqq\tfrac{1}{2}\|Ax-y\|^{2}$ under Gaussian noise), and $g$ is a regularization term that encodes prior beliefs about the class of admissible solutions. For instance, the $\ell^{1}$-norm ($g(x)\propto\lVert x\rVert_{1}$) promotes sparsity and underlies the theory of compressed sensing [9, 17, 6]. Similarly, total variation (TV), defined as $g(x)\propto\operatorname{TV}(x)\coloneqq\sum_{k}|x_{k}-x_{k-1}|$ favors piecewise-constant solutions and is widely used in image denoising [52, 10, 11, 36, 44]. More recently, much of research in imaging inverse problems has shifted towards learned regularizers [29, 18, 19, 24, 48, 22, 43, 33]. Without regularization, the problem is often ill-posed in the sense of Hadamard [21, 28]: Solutions may be non-unique, unstable, or nonexistent. Since the data-fidelity term is typically well-behaved, the regularizer largely determines the structure of solutions. More broadly, the regularizer largely dictates how optimization algorithms behave, so understanding its structure is crucial.
I-A Proximal Gradient Descent.
Gradient descent is a classical method for minimizing smooth (Lipschitz gradient) objective functions, with iterations
$x_{k+1}\leftarrow x_{k}-\gamma\nabla F(x_{k}).$ (4)
With appropriately chosen step size $\gamma>0$, the iterates $x_{k}$ converge to a stationary point [32]. However, gradient descent is not well-suited for nonsmooth objectives.
Proximal gradient descent (PGD) addresses this limitation by splitting the objective function as in (3). The data-fidelity term $f$ is typically smooth while the regularizer $g$ is often not (such as $\ell^{1}$ or $\operatorname{TV}$). At each iteration, PGD takes a gradient step on the data-fidelity term followed by a proximal step for the regularizer, yielding iterations of the form
[…]
Here $\operatorname{prox}_{\gamma g}$ denotes the proximal operator of $g$ (see Section II), originally introduced by Moreau in [40]. Convergence of PGD has been widely studied in both convex and non-convex settings and can be established under various assumptions on $f$ and $g$ with appropriately chosen step sizes $\gamma>0$ [35, Table 1].
I-B Plug-and-Play Methods
The plug-and-play (PnP) framework [58, 33] extends PGD by replacing the proximal step $\operatorname{prox}_{\gamma g}$ with a generic denoiser $D_{\sigma}$ [7]. This substitution is motivated by the interpretation of $\operatorname{prox}_{\gamma g}$ as the maximum a posteriori (MAP) solution to the denoising problem (2) under prior $p_{X}\propto e^{-\gamma g}$ [14, Eq. (10.13)], the exact problem denoisers are designed to solve. The PnP iteration is given by
[…]
and has been shown to yield state-of-the-art results when combined with powerful image denoisers such as BM3D [16] or a deep neural network (DNN) [31, 60]. Often, we use data fidelity term $f(x)\coloneqq\tfrac{1}{2}\|Ax-y\|^{2}$, in which case the PnP iteration becomes
[…]
A central challenge in the PnP framework is that while a proximal operator is always tied to a well-defined regularizer, many of the most effective denoising methods are not. This substitution therefore breaks the direct connection to an explicit optimization problem, making theoretical interpretations and convergence claims difficult. Much of the literature imposes strict assumptions on the denoiser to obtain convergence guarantees. Many of the most effective denoisers do not satisfy these assumptions, yet still achieve excellent results (for example [1]).
I-C Related Works.
When $f$ and $g$ are convex with $L$-smooth $f$, it is well known that PGD converges to a fixed point for appropriate step size $\gamma$ [45, Section 4.2]. This result can be proven via the machinery of monotone operator theory, the goal being to show that the composite PGD step (5) constitutes an averaged operator, guaranteeing convergence [15, 13, 3]. [^1] When $g$ is proper, lower-semicontinuous, convex, $\partial g$ is maximally monotone, making $\operatorname{prox}_{\gamma g}=(I+\gamma\partial g)^{-1}$ firmly-nonexpansive, i.e., $1/2$-averaged. The gradient step is also averaged for $L$-smooth $f$, making their composition an averaged operator.
In particular, convergence of PnP-PGD has been shown under tight assumptions on the denoiser, such as averaged/nonexpansive behavior [56] or suitable Lipschitz (e.g., residual-Lipschitz) bounds [53]. Many works impose these assumptions through architectural constraints on the learned denoisers via, e.g., spectral normalization [53] or through explicit training techniques [57, 8]. In [47], the authors propose a technique for learning a maximally monotone operator. Convergence of PnP has also been shown for specific denoisers, such as kernel denoisers [23], denoisers that can be written as the gradient step of a smooth function [30], bounded denoisers [12], and linear denoisers [42].
Instead of treating the denoiser as a black box with certain conditions, a complementary line of work leverages its intrinsic statistical structure. The MMSE estimator under Gaussian noise is a notable example. With the Tweedie/Stein identities [20, 55], the MMSE denoiser can be tied to the score function, endowing it with useful structural properties [38]. Further, [25] and [27] showed that the MMSE denoiser is $C^{\infty}$ and can be written as the proximal map of a regularizer (which is $C^{\infty}$) under various forward and noise models, and derived an explicit formula for this regularizer (28). This progress enabled the derivation of asymptotic convergence for both the ADMM [46] and PGD [59] variants of PnP using an MMSE denoiser. However, a gap in the theory remains: The explicit formula offers little insight into the behavior of the induced regularizer, and there are no nonasymptotic convergence rates tailored to MMSE denoisers.
I-D Contributions.
We make the following contributions.
  • Characterization of the implicit regularizer of the MMSE denoiser. We show in Theorem III.2 that the MMSE denoiser is the proximal map of a regularizer that admits an explicit form as an upper Moreau envelope, i.e., $\phi_{\mathrm{MMSE}}(x)$ is, up to a constant,
    $\sigma^{2}\sup_{z}\mathopen{}\mathclose{{\left(f_{Z}(z)-\tfrac{1}{2\sigma^{2}}\|z-x\|^{2}}}\right)$ (8)
    where $f_{Z}=-\log p_{Z}$ is the negative log marginal distribution in the noise model Eq. 2. This is an explicit and interpretable description of the regularizer implicit in the MMSE denoiser, and we believe that this identity is new.
  • Nonasymptotic convergence of MMSE-PGD via weak convexity. In Section III-A, we prove that $x\mapsto\phi_{\mathrm{MMSE}}(x)+\tfrac{1}{2}\|x\|^{2}$ is convex, hence $\phi_{\mathrm{MMSE}}$ is $1$-weakly convex. With this property we establish (to the best of our knowledge) the first nonasymptotic convergence guarantees for PGD-PnP with an MMSE denoiser in Theorem III.6. In particular, we show that, under an $L$-smooth data-fidelity term $f$ with $L<1$, PGD reaches an $\mathcal{O}(1/\sqrt{k})$-stationary point after $k$ iterations.
  • Experiments. In Section IV, we consider the following experimental setups: (i) A one-dimensional synthetic example validating the upper Moreau envelope form of the regularizer implicit in the MMSE denoiser by comparing a learned DNN estimator to its explicit form, (ii) Gaussian deblurring on the MNIST dataset [34], and (iii) computed tomography on the MayoCT dataset [37]. In both imaging tasks we observe the predicted sublinear $\mathcal{O}(1/\sqrt{k}$) convergence in the subgradient residual.
II Preliminaries
In this section, we collect some definitions and results used in the remainder of the paper. First, we recall the definition of the convex conjugate, also known as the Legendre-Fenchel transform or Fenchel conjugate. See for example [51].
Definition II.1.
For a function $f:\mathbb{R}^{n}\to\mathbb{R}$, the convex conjugate $f^{*}$ is given by
[…]
Definition II.2 (Lower Moreau Envelope).
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a function with $\gamma>0$. The lower Moreau envelope of $f$ is given by
$M_{\gamma}f(x)\coloneqq\inf_{y}\mathopen{}\mathclose{{\left(f(y)+\tfrac{1}{2\gamma}\|y-x\|^{2}}}\right).$ (10)
Definition II.3 (Upper Moreau Envelope).
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a function with $\gamma>0$. The upper Moreau envelope of $f$ is given by
$M^{\gamma}f(x)\coloneqq\sup_{y}\mathopen{}\mathclose{{\left(f(y)-\tfrac{1}{2\gamma}\|y-x\|^{2}}}\right).$ (11)
Note that the lower Moreau envelope is the usual Moreau envelope, but we make the distinction in this paper since we use both lower and upper Moreau envelopes. For the infimum defining the lower Moreau envelope to be achieved, $f$ must be lower-semicontinuous. An important property of the lower Moreau envelope $M_{\gamma}f$ is that it shares global minimizers with $f$ whenever $f$ is lower-semicontinuous, i.e. $\operatorname*{arg\,min}_{x}f(x)=\operatorname*{arg\,min}_{x}M_{\gamma}f(x).$ See [4] for further background.
Definition II.4 (Weak convexity).
Let $f:\mathbb{R}^{n}\to\mathbb{R}$, $\gamma>0$. $f$ is called $\gamma$-weakly convex if the mapping $x\mapsto f(x)+\tfrac{\gamma}{2}\|x\|^{2}$ is convex.
Crucial to our analysis is the following lemma, where we include a short proof for completeness since the proof in the original reference had typos.
Lemma II.5 ([4, Lemma 3]).
If $f:\mathbb{R}^{n}\to\mathbb{R}$ is locally bounded, $M^{\gamma}M_{\gamma}f(x)\leq f(x)$, and equality holds if and only if $f$ is $\tfrac{1}{\gamma}$-weakly convex.
Proof.
From the definition,
$\phantom{{}={}}M^{\gamma}M_{\gamma}f(x)$
$=\sup_{y}\inf_{z}\mathopen{}\mathclose{{\left(f(z)+\tfrac{1}{2\gamma}\|z-y\|^{2}-\tfrac{1}{2\gamma}\|y-x\|^{2}}}\right)$
$\leq\sup_{y}f(x)=f(x),$ (12)
by taking $z=x$. We now write
$\phantom{{}={}}M^{\gamma}M_{\gamma}f(x)+\tfrac{1}{2\gamma}\|x\|^{2}$
[…]
The convex conjugate $g^{*}$ evaluated at $y/\gamma$ can be written as
$g^{*}(y/\gamma)$ $=\sup_{z}\mathopen{}\mathclose{{\left(\langle y/\gamma,z\rangle-g(z)}}\right)$
$=-\inf_{z}\mathopen{}\mathclose{{\left(g(z)-\langle y/\gamma,z\rangle}}\right),$ (14)
so that (13) can be written as
$\phantom{{}={}}\sup_{y}\mathopen{}\mathclose{{\left(-g^{*}(y/\gamma)+\langle y/\gamma,x\rangle}}\right)$
$=\sup_{y}\mathopen{}\mathclose{{\left(\langle y,x\rangle-g^{*}(y)}}\right)=g^{**}(x).$ (15)
Therefore, we have
$M^{\gamma}M_{\gamma}f(x)+\tfrac{1}{2\gamma}\|x\|^{2}=\mathopen{}\mathclose{{\left(f+\tfrac{1}{2\gamma}\|\cdot\|^{2}}}\right)^{**}(x),$ (16)
and the result follows when we recall that a function equals its biconjugate if and only if it is convex [50, Section 12.2]. ∎
We now turn to the proximal operator, a central tool in optimization. We will later show that MMSE estimators themselves admit a proximal representation.
Definition II.6 (Proximal operator).
Let $\gamma>0$, and $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ be proper and lower-semicontinuous. The proximal operator of $f$ with parameter $\gamma$ is the mapping
[…]
Note that $\operatorname{prox}_{\gamma f}$ is often set-valued. By coercivity of the quadratic term, $\operatorname{prox}_{\gamma f}(x)$ is non-empty for every $x\in\mathbb{R}^{n}$. For a proper, lower-semicontinuous function $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$, the infimum defining the lower Moreau envelope is attained at every minimizer of the proximal problem; that is,
[…]
for all $x\in\mathbb{R}^{n}$. When $f$ is a convex, lower-semicontinuous proper function, it is well known that $\operatorname{prox}_{\gamma f}$ and $M_{\gamma}f$ are also related by the following formula due to Moreau [41]:
$\nabla M_{\gamma}f(x)=\tfrac{1}{\gamma}(x-\operatorname{prox}_{\gamma f}(x)).$ (19)
Although the proximal operator of a non-convex function is generally set-valued, when $\operatorname{prox}_{\gamma f}$ is an MMSE denoiser, it is always single-valued and continuous, which is the setting of our main results. This motivates the following extension of the Moreau gradient identity.
Theorem II.7 (Extension of the Moreau Gradient Identity).
Let $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ be proper and lower-semicontinuous. If $\operatorname{prox}_{\gamma f}(x)$ is single-valued and continuous for all $x$, then the gradient of the lower Moreau envelope exists and takes the form
$\nabla M_{\gamma}f(x)=\tfrac{1}{\gamma}(x-\operatorname{prox}_{\gamma f}(x)),\quad x\in\mathbb{R}^{n}.$ (20)
Proof.
Let $h(x,z)\coloneqq f(z)+\tfrac{1}{2\gamma}\|z-x\|^{2}$. We have for all $z\in\mathbb{R}^{n}$, $t\in\mathbb{R}$,
[…]
From the definition, $M_{\gamma}f(x)=\min_{z}h(x,z)$. Fix some $\bar{x}\in\mathbb{R}^{n}$ where there is a unique minimizer $\{\bar{z}\}=\operatorname{prox}_{\gamma f}(\bar{x})$ by assumption. Take some $u\in\mathbb{R}^{n}$. There is some $\{z_{t}\}=\operatorname{prox}_{\gamma f}(\bar{x}+tu)$ so that
$\phantom{{}={}}M_{\gamma}f(\bar{x}+tu)=h(\bar{x}+tu,z_{t})$
$=h(\bar{x},z_{t})+\tfrac{t}{\gamma}\langle\bar{x}-z_{t},u\rangle+\tfrac{t^{2}}{2\gamma}\|u\|^{2}$
$\geq M_{\gamma}f(\bar{x})+\tfrac{t}{\gamma}\langle\bar{x}-z_{t},u\rangle+\tfrac{t^{2}}{2\gamma}\|u\|^{2}.$ (22)
Subtracting and dividing through by $t$, we get
$\phantom{{}={}}\liminf_{t\to 0}\tfrac{M_{\gamma}f(\bar{x}+tu)-M_{\gamma}f(\bar{x})}{t}$
$\geq\liminf_{t\to 0}\mathopen{}\mathclose{{\left(\tfrac{1}{\gamma}\langle\bar{x}-z_{t},u\rangle+\tfrac{t}{2\gamma}\|u\|^{2}}}\right)$
$=\tfrac{1}{\gamma}\langle\bar{x}-\bar{z},u\rangle,$ (23)
where the last equality is by the assumption that $\operatorname{prox}_{\gamma f}$ is continuous. Next,
$\phantom{{}={}}M_{\gamma}f(\bar{x}+tu)\leq h(\bar{x}+tu,\bar{z})$
$=M_{\gamma}f(\bar{x})+\tfrac{t}{\gamma}\langle\bar{x}-\bar{z},u\rangle+\tfrac{t^{2}}{2\gamma}\|u\|^{2}.$ (24)
Again, subtracting and dividing through by $t$,
$\phantom{{}={}}\limsup_{t\to 0}\tfrac{M_{\gamma}f(\bar{x}+tu)-M_{\gamma}f(\bar{x})}{t}$
$\leq\limsup_{t\to 0}\mathopen{}\mathclose{{\left(\tfrac{1}{\gamma}\langle\bar{x}-\bar{z},u\rangle+\tfrac{t}{2\gamma}\|u\|^{2}}}\right)$
$=\tfrac{1}{\gamma}\langle\bar{x}-\bar{z},u\rangle.$ (25)
Combining (II) and (II), we have that the directional derivative of $M_{\gamma}f$ at $\bar{x}$ in direction $u$ satisfies $\nabla_{u}M_{\gamma}f(\bar{x})=\tfrac{1}{\gamma}\langle\bar{x}-\bar{z},u\rangle$. Therefore, the gradient is
$\nabla M_{\gamma}f(x)=\tfrac{1}{\gamma}(\bar{x}-\bar{z})=\tfrac{1}{\gamma}\mathopen{}\mathclose{{\left(\bar{x}-\operatorname{prox}_{\gamma f}(\bar{x})}}\right).$ (26)
Many similar versions of this identity are well known in the literature. To our knowledge, this exact formulation has not been explicitly stated.
We next review key results on MMSE denoisers and state their connection to proximal mappings, beginning with Tweedie’s formula.
Theorem II.8 (Tweedie’s Formula).
Let $X\sim p_{X}$, $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ be independent random variables in $\mathbb{R}^{n}$, and let $Z=X+\epsilon$, $\psi_{\sigma}(z)\coloneqq\mathbb{E}[X|Z=z]$. Then,
[…]
where $p_{Z}=p_{X}*\mathcal{N}(0,\sigma^{2}I)$.
For a thorough introduction to Tweedie’s formula, we refer the reader to [20]. Note that Tweedie’s Formula makes no assumptions on the prior density $p_{X}$, which may even include discrete atoms.
Theorem II.9 ([27] and [25, Corollary 1]).
Let $X\sim p_{X}$ and $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ be independent random variables in $\mathbb{R}^{n}$, and let $Z=X+\epsilon$. Assume $p_{X}$ is non-degenerate.222i.e., there is no pair $v\in\mathbb{R}^{n}$, $c\in\mathbb{R}$ such that $\langle X,v\rangle=c$, almost surely. The MMSE estimator under (2) given by $\psi_{\sigma}(z)\coloneqq\mathbb{E}[X|Z=z]$ has the following properties:
  1. 1.
    It is one-to-one from $\mathbb{R}^{n}$ onto its image.
  2. 2.
    It is $C^{\infty}$, and so is its inverse.
  3. 3.
    It can be written as $\psi_{\sigma}(z)=\operatorname{prox}_{\phi_{\mathrm{MMSE}}}(z)$, where
    $\phi_{\mathrm{MMSE}}(z)\coloneqq$
[…]
    where $f_{Z}=-\log p_{Z}$.
  4. 4.
    $\phi_{\mathrm{MMSE}}$ is $C^{\infty}$.
III Main Results
In this section, we present our theoretical contributions. We begin by introducing a new structural representation of the implicit regularizer underlying the MMSE denoiser. This characterization provides greater interpretability and serves as the foundation for deriving novel nonasymptotic convergence guarantees.
III-A A New Characterization of the MMSE Regularizer.
It is known (Theorem II.9) that the MMSE denoiser is the proximal map of an infinitely differentiable penalty. Yet beyond this, the structure of the associated penalty $\phi_{\mathrm{MMSE}}$ remains unclear. The only known explicit expression of $\phi_{\mathrm{MMSE}}$ (28) is cumbersome and offers little analytical insight. In this section, we show that $\phi_{\mathrm{MMSE}}$ admits an upper Moreau envelope representation in terms of $f_{Z}$, yielding useful insights, namely $1$-weak convexity, which we will leverage to prove stronger convergence results for the PGD using an MMSE denoiser.
For the following, we will assume the setting of the denoising problem (2), letting $X\sim p_{X}$, $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ be independent random variables in $\mathbb{R}^{n}$, $Z=X+\epsilon$, so that $p_{Z}=p_{X}*\mathcal{N}(0,\sigma^{2}I)$. We will let $f_{Z}\coloneqq-\log p_{Z}$.
Lemma III.1.
The negative log marginal distribution from the denoising problem (2) can be written as
$f_{Z}=\tfrac{1}{\sigma^{2}}M_{1}\phi_{\mathrm{MMSE}}+C_{x_{0}},$ (29)
where $\phi_{\mathrm{MMSE}}$ is as in (28) and $C_{x_{0}}$ takes the form
[…]
for any choice of $x_{0}\in\mathbb{R}^{n}$.
Proof.
Pick $x\in\mathbb{R}^{n}$. By Theorem II.8,
$x-\operatorname{prox}_{\phi_{\mathrm{MMSE}}}(x)$ $=-\sigma^{2}\nabla\log p_{Z}(x)$
$=\sigma^{2}\nabla f_{Z}(x),$ (31)
which gives us
$\sigma^{2}\nabla f_{Z}(x)=\nabla M_{1}\phi_{\mathrm{MMSE}}(x),$ (32)
when combined with Theorem II.7. Integrating gives the result. ∎
Theorem III.2.
The regularizer implicit in the MMSE denoiser can be written as
[…]
for all $x$ in the image of $\psi_{\sigma}$.
Proof.
Pick some $x\in\mathbb{R}^{n}$. By Section II,
$\phi_{\mathrm{MMSE}}(x)\geq M^{1}M_{1}\phi_{\mathrm{MMSE}}(x),$ (34)
which, by the previous result, we can rewrite as
$\phi_{\mathrm{MMSE}}(x)$ $\geq M^{1}(\sigma^{2}f_{Z}-\sigma^{2}C_{x_{0}})(x)$
$=\sigma^{2}M^{\sigma^{2}}f_{Z}(x)-\sigma^{2}C_{x_{0}}.$ (35)
For the reverse inequality, take some $x$ in the image of $\psi_{\sigma}$. There exists some $z\in\mathbb{R}^{n}$ so that $x=\psi_{\sigma}(z)=\operatorname{prox}_{\phi_{\mathrm{MMSE}}}(z)$. The infimum defining $M_{1}\phi_{\mathrm{MMSE}}$ is achieved at $x$ (see (18)), so that
$\phi_{\mathrm{MMSE}}(x)+\tfrac{1}{2}\|x-z\|^{2}$ $=M_{1}\phi_{\mathrm{MMSE}}(z)$
$=\sigma^{2}f_{Z}(z)-\sigma^{2}C_{x_{0}},$ (36)
where the second line holds by Section III-A. Rearranging, we have
$\phi_{\mathrm{MMSE}}(x)$ $=\sigma^{2}\mathopen{}\mathclose{{\left(f_{Z}(z)-\tfrac{1}{2\sigma^{2}}\|x-z\|^{2}}}\right)-\sigma^{2}C_{x_{0}}$
$\leq\sigma^{2}M^{\sigma^{2}}f_{Z}(x)-\sigma^{2}C_{x_{0}}.$ (37)
We illustrate this result in Fig. 1 and Fig. 2 under a variety of mixture of Gaussian and mixture of Laplacian priors. An overview of the experiment is provided in Section IV-A.
Figure 1: Calculated and learned regularizers for an MMSE denoiser under mixture-of-Gaussian priors with unit Gaussian noise. The calculated regularizer is derived via Theorem III.2.
Figure 2: Calculated and learned regularizers for an MMSE denoiser under mixture-of-Laplacian priors with unit Gaussian noise. The calculated regularizer is derived via Theorem III.2.
Corollary III.3.
The regularizer $\phi_{\mathrm{MMSE}}$ implicit in the MMSE denoiser is $1$-weakly convex on the image of $\psi_{\sigma}$.
Proof.
From Theorem III.2,
$\phi_{\mathrm{MMSE}}$ $=\sigma^{2}M^{\sigma^{2}}f_{Z}-\sigma^{2}C_{x_{0}}$
$=M^{1}(\sigma^{2}(f_{Z}-C_{x_{0}}))=M^{1}M_{1}\phi_{\mathrm{MMSE}},$ (38)
by Section III-A. This holds only if $\phi_{\mathrm{MMSE}}$ is 1-weakly convex by Section II. ∎
We now leverage this 1-weak convexity to derive nonasymptotic guarantees for PnP-PGD.
III-B Proximal Gradient Descent Convergence Results.
When the objective $F$ is smooth, the map $x\mapsto\nabla F(x)$ is continuous. Therefore, measuring convergence via the gradient norm $\|\nabla F(x_{k})\|$ is natural; small gradient norms certify proximity to a stationary point. In the non-smooth, non-convex setting, the gradient may not exist and is generalized to the subgradient [51, Chapter 8].
Definition III.4 (Subgradient).
For a function $F:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$, the subgradient $\partial F(x)$ is given by
$\partial F(x)\coloneqq\mathopen{}\mathclose{{\left\{v\in\mathbb{R}^{n}\;\middle|\;\begin{aligned} &F(y)\geq F(x)+\langle v,y-x\rangle,\\ &\text{for all }y\in\mathbb{R}^{n}\end{aligned}}}\right\}$ (39)
Recall that $x^{\star}$ is a stationary point for $F$ if and only if $0\in\partial F(x^{\star})$. This motivates measuring progress by approximate stationarity. In the nonsmooth, non-convex setting, the natural goal is to find an $\varepsilon$-stationary point, i.e. an $x$ satisfying
$\operatorname{dist}(0,\partial F(x))\leq\varepsilon,$ (40)
where
[…]
In non-convex, nonsmooth problems, the stationary residual (41) may oscillate, even while it trends downward. This can depend on the forward operator as well as the data. To account for this behavior, nonasymptotic convergence results are typically stated in terms of the best iterate so far. See for example [2]. It is standard to state that an algorithm reaches an $\varepsilon$-stationary point in the first $k$ iterations if
$\min_{1\leq j\leq k}\operatorname{dist}(0,\partial F(x_{j}))\leq\varepsilon.$ (42)
For our investigation, we prove the following intermediary result, which is a simple improvement over [5, Theorem 5.1]. This result will be used to prove convergence of PGD using an MMSE denoiser.
Lemma III.5.
Let $\{x_{k}\}$ be iterates of proximal gradient descent defined in (5), where $g$ is proper, lower-semicontinuous, $\rho$-weakly convex, and $\nabla f$ is $L$-Lipschitz continuous. Suppose $\gamma\in\mathopen{}\mathclose{{\left(0,\tfrac{2}{L+\rho}}}\right)$. Then, for all $k\geq 1$,
$\min_{2\leq j\leq k+1}\operatorname{dist}(0,\partial F(x_{j}))\leq k^{-1/2}C\sqrt{F(x_{1})-F^{*}},$ (43)
where $F$ is as in (3),
[…]
and
$C=\tfrac{\tfrac{1}{\gamma}+L}{\sqrt{\tfrac{1}{\gamma}-\tfrac{L+\rho}{2}}}.$ (45)
In other words, there is an $\epsilon$-stationary point in the first $k+1$ iterates with $\epsilon\sim\mathcal{O}(1/\sqrt{k})$.
Proof.
We first note that $\widetilde{g}(x)\coloneqq g(x)+\tfrac{\rho}{2}\|x\|^{2}$ is convex, so
[…]
for any $s\in\partial\widetilde{g}(x)$. For iteration $x_{k+1}$, let $s_{k+1}\in\partial g(x_{k+1})$, so that $s_{k+1}+\rho x_{k+1}\in\partial\widetilde{g}(x_{k+1})$. We then have
$g(x_{k})+\tfrac{\rho}{2}\|x_{k}\|^{2}$ $\geq g(x_{k+1})+\tfrac{\rho}{2}\|x_{k+1}\|^{2}$
$+\langle s_{k+1}+\rho x_{k+1},x_{k}-x_{k+1}\rangle,$ (47)
using Eq. 46 with $y=x_{k}$, $x=x_{k+1}$, and $s=s_{k+1}+\rho x_{k+1}$. We will write this as
[…]
$-\tfrac{\rho}{2}\|x_{k}-x_{k+1}\|^{2}.$ (48)
$\nabla f$ is $L$-Lipschitz, so
[…]
$+\tfrac{L}{2}\|x_{k+1}-x_{k}\|^{2}.$ (49)
Setting $s_{k+1}\coloneq\tfrac{1}{\gamma}(x_{k}-x_{k+1})-\nabla f(x_{k})\in\partial g(x_{k+1})$ and adding (III-B) and (III-B),
[…]
Next, let $w_{k+1}\coloneqq\tfrac{1}{\gamma}(x_{k+1}-x_{k})+\nabla f(x_{k+1})-\nabla f(x_{k})$, an element of $\partial(f+g)(x_{k+1})$,
$\|w_{k+1}\|^{2}$ $\leq\mathopen{}\mathclose{{\left(\tfrac{1}{\gamma}+L}}\right)^{2}\|x_{k+1}-x_{k}\|^{2}$
$\leq C^{2}\mathopen{}\mathclose{{\left(F(x_{k})-F(x_{k+1})}}\right),$ (51)
by (50), where $C=\tfrac{\tfrac{1}{\gamma}+L}{\sqrt{\tfrac{1}{\gamma}-\tfrac{L+\rho}{2}}}$. Next, we have a simple telescoping sum,
$\min_{1\leq j\leq k}\|w_{j+1}\|^{2}$ $\leq\tfrac{1}{k}\sum_{j=1}^{k+1}\|w_{j}\|^{2}$
$\leq\tfrac{1}{k}C^{2}(F(x_{1})-F(x_{k+1}))$
$\leq C^{2}(F(x_{1})-F^{*}),$ (52)
using the definition in (44). Taking a square root and recalling that $\operatorname{dist}(0,\partial F(x_{j+1}))\leq\|w_{j+1}\|$ gives us the result. ∎
Convergence of PGD using an MMSE denoiser has been previously established for a range of step sizes $\gamma>0$. In particular, [59] showed that PnP-MMSE converges for any $\gamma\leq 1/L$, but the objective it converges to is not the objective of interest, i.e.,
[…]
but instead it converges to a scaled objective of the form
$F(x)=f(x)+\tfrac{1}{\gamma}\phi_{\mathrm{MMSE}}(x).$ (54)
Furthermore, their convergence is only asymptotic. On the other hand, in this paper, we establish a nonasymptotic convergence result for the objective (53). Our result is summarized in the following theorem.
Theorem III.6.
Consider the PnP iteration (6) using $D_{\sigma}\coloneqq\psi_{\sigma}$, the MMSE denoiser with implicit step size $\gamma=1$. If $\nabla f$ is $L$-Lipschitz continuous with $L<1$, there exists an $\mathcal{O}(1/\sqrt{k})$-stationary point among $\{x_{j}\}_{j=1}^{k+1}$.
Proof.
Apply Section III-B with $\rho=1$, since $\phi_{\mathrm{MMSE}}$ is $1$-weakly convex. Note that the iterates $x_{j}$ lie in the image of $\psi_{\sigma}$ for $j\geq 2$, so Theorem III.2 applies to them. ∎
IV Experiments
IV-A Illustration of Theorem III.2.
Figure 3: MMSE denoiser, convex potential, and implicit MMSE regularizer for a Laplacian prior $p_{X}(x)=e^{-|x|}$, plotted against reference curves. LPN denotes the learned proximal network. This explicates what is shown in [22, Figure 2], but plotted on a wider axis to show the “curling down” behavior of LPNs.
We consider the denoising problem (2) under mixture-of-Gaussian and mixture-of-Laplacian priors with unit Gaussian noise. We train a gradient-output softplus input convex neural network (ICNN) to approximate the MMSE denoiser and calculate its implicit prior via [26, Theorem 3]. We compare this with the implicit regularizer calculated using the explicit upper Moreau envelope formula from Theorem III.2. Figures 1 and 2 compare the calculated $\phi_{\mathrm{MMSE}}$ and the learned estimate, along with reference curves $f_{X}\coloneqq-\log p_{X}$ and $f_{Z}\coloneqq-\log p_{Z}$. The learned and calculated regularizers are nearly identical where the samples are most concentrated.
We observe that the learned implicit regularizer “curls downward” outside the data’s central region. This behavior is intrinsic to ReLU-based neural networks (and their softplus approximations). It is well known that feed-forward networks with piecewise-affine activations yield piecewise-affine mappings [49, 39]. In particular, ReLU is piecewise-affine and is the uniform limit of the softplus family as $\beta\to\infty$. Thus, an ICNN with softplus activations is a smooth approximation of a ReLU-ICNN and is therefore “nearly” piecewise-affine. Outside the support of the training data, the learned convex potential becomes nearly affine, and its gradient (the denoiser) saturates at an approximately constant value. Consequently, the implicit regularizer appears to bend downward outside the central region. We emphasize that this behavior arises as a result of the ReLU-approximating architecture, not from the MMSE loss or our implementation. To illustrate this, in Figure 3 we reproduce the setup of [22, Fig. 2] and simply extend the axis range. The extended view makes the affine-tail/saturation effect apparent.
IV-B Gaussian Blurring.
We next consider PGD for the Gaussian blurring inverse problem (1), where $X$ is sampled from the MNIST data set of handwritten digits [34] (resolution of $28\times 28$). $A$ is a Gaussian blurring operator, computed by convolution with a $3\times 3$ Gaussian kernel with $\sigma^{2}=1$, and measurement noise $\epsilon$ being a zero-mean standard Gaussian with variance $\sigma^{2}=0.04$. The Gaussian kernel has unit mass, giving the blur an operator norm $\lVert A\rVert_{\mathrm{op}}=1$.
We use data-fidelity term $f(x)\coloneqq\tfrac{1}{2}\|Ax-y\|^{2}$ normalized to have $L<1$. We use an MMSE denoiser trained with noise $\sigma^{2}=0.04$ for the proximal step. We confirm the efficacy of this algorithm on 5 test examples in Figure 4. At this noise level, the observed blurry images have a PSNR of $\approx 15$ dB, while the reconstructions reach PSNR of $\approx 20$ dB.
In Figure 5, we plot $\operatorname{dist}(0,{\partial}F(x_{k}))$ over $50$ iterations for the first test image in Figure 4, showing clear sublinear fixed-point convergence, as predicted by Theorem III.6.
Figure 4: PNP-PGD results on the MNIST dataset under Gaussian blur.
Figure 5: Stationary residual (41) over 50 iterations for PNP-PGD using an MMSE denoiser on the MNIST dataset under Gaussian blur.
IV-C Computed Tomography.
We study the CT inverse problem with forward operator $A$ given by the discretized Radon transform, normalized to $\|A\|_{\mathrm{op}}<1$. Training images are random $128\times 128$ patches sampled from the MayoCT dataset [37] (resolution of $512\times 512$).
We test the PGD algorithm using the MMSE denoiser on randomly selected regions of the Shepp–Logan phantom [54], which captures a wide range of structures relevant in computed tomography. We corrupt measurements with zero-mean Gaussian noise to yield input PSNR levels of $20$, $24$, and $28$ dB, ranges commonly studied for low-dose CT (LDCT). See for example [61, Table 6]. After $5$ iterations of PGD we obtain reconstructions in the range of 37373838 dB which we display in Figs. 6 and 7.
Figure 6: Randomly selected $128\times 128$ region of the Shepp–Logan phantom. Clean, noisy/blurred, and PnP-PGD reconstruction with an MMSE denoiser after 5 iterations. Gaussian noise; per-panel PSNR/SSIM are denoted.
Figure 7: Randomly selected $128\times 128$ region of the Shepp–Logan phantom. Clean, noisy/blurred, and PnP-PGD reconstruction with an MMSE denoiser after 5 iterations. Gaussian noise; per-panel PSNR/SSIM are denoted.
In Fig. 8, we plot the stationary residual (41) for 160 iterations and observe behavior consistent with our findings in Theorem III.6
Figure 8: Stationary residual (41) over 160 iterations for PNP-PGD on the MayoCT dataset using an MMSE denoiser.
IV-D Architectural Details.
We build upon the implementation of the Input Convex Neural Network (ICNN) in [22] to approximate the MMSE denoiser. The authors produced excellent results with softplus activations and a mixture of linear and convolutional layers. [^3] The exact PyTorch implementation is available at https://github.com/ZhenghanFang/learned-proximal-networks The network uses layers of the form
$h_{0}$ $=\sigma(A_{0}x+b_{0}),$ (55)
$h_{k+1}$ $=\sigma(W_{k+1}h_{k}+A_{k+1}x+b_{k+1}),$ (56)
with hidden layers $W_{k}\in\mathbb{R}^{H\times H}$, skip connections $A_{k}\in\mathbb{R}^{H\times H}$, and softplus activation $\sigma(x)=\tfrac{1}{\beta}\log\mathopen{}\mathclose{{\left(1+\exp(\beta x)}}\right)$ applied element-wise ($\beta=10$). Softplus is a differentiable approximation of ReLU, so that the denoiser can be defined as the gradient $\psi_{\theta}\coloneqq\nabla g_{\theta}$ of potential $g_{\theta}:\mathbb{R}^{n}\to\mathbb{R}$ given by
$g_{\theta}(x)\coloneqq w_{\mathrm{out}}^{T}h_{L}+a_{\mathrm{out}}^{T}x+b_{\mathrm{out}}.$ (57)
To ensure convexity in $x$, all hidden layers $W_{k}$ are clipped so that they have only non-negative entries. Note that a convolutional layer can be represented as a matrix. For the MayoCT and MNIST experiments, the hidden layers are clipped in this form, not in kernel form. Because the activation is convex and non-decreasing, these constraints make $g_{\theta}$ convex, guaranteeing that its gradient ($\psi_{\theta}$) is a proximal operator [26, Corollary 1]. Note that no restrictions on the skip-connections or biases are necessary.
IV-E Training Details.
For the illustration of Theorem III.2, the data was divided into batches of $4,000$ with $2,000$ iterations at learning rate $10^{-3}$. For Gaussian deblurring, the dataset was divided into batches of 500 with noise of $\sigma^{2}=0.04$. The training schedule is detailed in Table I.
TABLE I: MNIST and MayoCT Denoiser Learning-rate Schedule
Phase Learning rate Steps
1 $1\times 10^{-3}$ 10,000
2 $3\times 10^{-4}$ 15,000
3 $1\times 10^{-4}$ 10,000
4 $3\times 10^{-5}$ 5,000
For computed tomography, the dataset was divided into batches of 64 with noise level of $\sigma^{2}=0.01$. The forward operator was implemented using the ODL [^4] Documentation available at https://odlgroup.github.io/odl/. Python library and normalized to have operator norm $\|A\|_{op}<1$. Training examples were identical to Table I. All training and testing was done on an Nvidia 4070 GPU using the Ubuntu 24.04 Windows Subsystem for Linux.
Remark IV.1.
We note that while the restriction of $L<1$ in Theorem III.6 may seem potentially restrictive, any $L$-smooth data-fidelity term can be scaled appropriately to meet this condition. For example, in our computed tomography experiment, the forward operator, a discretized Radon transform is $L$-smooth with $L\approx 450$. We simply scale the data fidelity term appropriately and still obtain strong empirical results with the desired convergence rate.
V Conclusion
We presented a novel representation of the MMSE denoiser’s implicit regularizer in Theorem III.2 and used it to derive nonasymptotic convergence results for PnP-PGD using an MMSE denoiser in Theorem III.6. Our analysis is strictly tied to the Gaussian-noise MMSE setting. Tweedie’s formula, for example, does not transfer to other noise models, so we suspect this flavor of result is unique to the Gaussian noise setting. However, we expect analogous nonasymptotic guarantees to be attainable for other PnP schemes that use MMSE denoisers under Gaussian noise, such as PnP-ADMM or the proximal point method. Exploration of these extensions constitutes future work toward more predictable PnP solvers.
Acknowledgment
The authors would like to thank Stanislas Ducotterd for helpful feedback on this work.