### v6

# Nonasymptotic Convergence Rates for Plug-and-Play Methods With MMSE Denoisers

Abstract
It is known that the minimum-mean-squared-error (MMSE) denoiser under Gaussian noise can be written as a proximal operator, which suffices for asymptotic convergence of plug-and-play (PnP) methods but does not reveal the structure of the induced regularizer or give convergence rates. We show that the MMSE denoiser corresponds to a regularizer that can be written explicitly as an upper Moreau envelope of the negative log-marginal density, which in turn implies that the regularizer is 1-weakly convex on its domain. Using this property, we derive (to the best of our knowledge) the first sublinear convergence guarantee for PnP proximal gradient descent with an MMSE denoiser. We validate the theory with a one-dimensional synthetic study that recovers the implicit regularizer. We also validate the theory with imaging experiments (deblurring and computed tomography), which exhibit the predicted sublinear behavior.
I Introduction
In this work, we focus on the setting of linear inverse problems, where $\overline{x}\in\mathbb{R}^{n}$ is the true signal, $y\in\mathbb{R}^{m}$ is the observed signal obtained by
[…]
for linear forward operator $A\in\mathbb{R}^{m\times n}$ and measurement noise $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$. The task of recovering $\overline{x}$ from $y$ is extensively studied in statistics and signal processing. A special case is the denoising problem, obtained when $A=I$:
[…]
Although simpler, denoising is directly useful in solving general inverse problems, often as an intermediate step.
Recovering $\overline{x}$ from $y$ is often posed as the minimization of a regularized objective function
[…]
where $f$ is a data-fidelity term (typically $f(x)\coloneqq\tfrac{1}{2}\|Ax-y\|^{2}$ under Gaussian noise), and $g$ is a regularization term that encodes prior beliefs about the class of admissible solutions. For instance, the $\ell^{1}$-norm ($g(x)\propto\lVert x\rVert_{1}$) promotes sparsity and underlies the theory of compressed sensing [11, 19, 8]. Similarly, total variation (TV), defined as $g(x)\propto\operatorname{TV}(x)\coloneqq\sum_{i}|x_{i}-x_{i-1}|$ for $1$-D signals, favors piecewise-constant solutions and is widely used in image denoising [57, 13, 12, 42, 51]. Without regularization, the problem is often ill-posed in the sense of Hadamard [23, 32]: Solutions may be non-unique, unstable, or nonexistent. Since the data-fidelity term is typically well-behaved, the regularizer largely determines the structure of solutions. More broadly, the regularizer largely dictates how optimization algorithms behave, so understanding its structure is crucial.
I-A Proximal Gradient Descent
Gradient descent is a classical method for minimizing smooth (Lipschitz gradient) objective functions, with iterations
$x_{k+1}=x_{k}-\gamma\nabla F(x_{k}).$ (4)
With appropriately chosen step size $\gamma>0$, the iterates $x_{k}$ converge to a stationary point [38]. However, gradient descent is not well-suited for nonsmooth objectives.
Proximal gradient descent (PGD) addresses this limitation by splitting the objective function as in (3). The data-fidelity term $f$ is typically smooth while the regularizer $g$ is often not (such as $\ell^{1}$ or $\operatorname{TV}$). At each iteration, PGD takes a gradient step on the data-fidelity term followed by a proximal step for the regularizer, yielding iterations of the form
[…]
Here $\operatorname{prox}_{\gamma g}$ denotes the proximal operator of $g$ (see Section II), originally introduced by Moreau in [46]. Convergence of PGD has been widely studied in both convex and nonconvex settings and can be established under various assumptions on $f$ and $g$ with appropriately chosen step sizes $\gamma>0$ [41, Table 1].
I-B Plug-and-Play Methods
The plug-and-play (PnP) framework [64, 39] extends PGD by replacing the proximal step $\operatorname{prox}_{\gamma g}$ with a generic denoiser $D_{\sigma}$ [9]. This substitution is motivated by the interpretation of $\operatorname{prox}_{\gamma g}$ as the maximum a posteriori (MAP) solution to the denoising problem (2) under prior $p_{X}\propto\mathrm{e}^{-\gamma g}$ [15, Eq. (10.13)], the exact problem denoisers are designed to solve. The PnP-PGD iteration is given by
[…]
and has been shown to yield state-of-the-art results when combined with powerful image denoisers such as BM3D [18] or a deep neural network (DNN) [37, 66]. Often, we use data fidelity term $f(x)\coloneqq\tfrac{1}{2}\|Ax-y\|^{2}$, in which case the PnP iteration becomes
[…]
A central challenge in the PnP framework is that while a proximal operator is always tied to a well-defined regularizer, many of the most effective denoising methods are not. This substitution therefore breaks the direct connection to an explicit optimization problem, making theoretical interpretations and convergence claims difficult. Much of the literature imposes strict assumptions on the denoiser to obtain convergence guarantees. Many of the most effective denoisers do not satisfy these assumptions, yet still achieve excellent results [2].
TABLE I: Summary of related convergence guarantees for plug-and-play proximal gradient descent (PnP-PGD). Note $L_{f}$ is the Lipschitz constant for $\nabla f$, $L_{g_{\sigma}}$ is the Lipschitz constant for $\nabla g_{\sigma}$, $\gamma$ is the step size, and $K$ is the iterate.
Work Assumptions on data-fidelity term $f$ Assumptions on denoiser $D_{\sigma}$ Convergence rate
Classical PGD [6] Convex, $L_{f}$-smooth $D_{\sigma}=\operatorname{prox}_{g}$, where $g$ is convex $\mathcal{O}(1/\sqrt{K})$
Lipschitz condition on the denoiser [58] $L_{f}$-smooth, $\mu$-Strongly convex $D_{\sigma}-I$ is $\varepsilon$-Lipschitz for $\varepsilon<\tfrac{2\mu}{L_{f}-\mu}$ Asymptotic
Gradient-step denoisers [34, 35, 36] $L_{f}$-smooth, $\gamma L_{f}<\tfrac{L_{g_{\sigma}}+2}{L_{g_{\sigma}}+1}$ $D_{\sigma}=I-\nabla g_{\sigma}$, where $L_{g_{\sigma}}<1$ $\mathcal{O}(1/\sqrt{K})$
Prior work with MMSE denoisers [65] $L_{f}$-smooth, $L_{f}<1$ MMSE denoiser Asymptotic
This paper (Section III-B) $L_{f}$-smooth, $L_{f}<1$ MMSE denoiser $\mathcal{O}(1/\sqrt{K})$
I-C Related Works
When $f$ and $g$ are convex with $L$-smooth $f$, it is well known that PGD converges to a fixed point for appropriate step size $\gamma$ [52, Section 4.2]. This result can be proven via the machinery of monotone operator theory, the goal being to show that the composite PGD step (5) constitutes an averaged operator, guaranteeing convergence [16, 17, 5]. [^1] When $g$ is proper, lower-semicontinuous, convex, $\partial g$ is maximally monotone, making $\operatorname{prox}_{\gamma g}=(I+\gamma\partial g)^{-1}$ firmly-nonexpansive, i.e., $1/2$-averaged. The gradient step is also averaged for $L$-smooth $f$, making their composition an averaged operator.
In particular, convergence of PnP-PGD has been shown under tight assumptions on the denoiser, such as nonexpansiveness [61] or suitable Lipschitz (e.g., residual-Lipschitz) bounds [58]. Many works impose these assumptions through architectural constraints on the learned denoisers via, e.g., spectral normalization [58] or through explicit training techniques [62, 10]. In [54], the authors propose a technique for learning a maximally monotone operator. Convergence of PnP has also been shown for specific denoisers, such as kernel denoisers [26, 3], gradient-step (GS) denoisers [36, 34, 35], bounded denoisers [14], and linear denoisers [48, 25]. A recent line of work [36] has significantly relaxed the assumptions on GS denoisers. We note that this setting (see [36, Proposition 3.1]) still does not apply to the MMSE denoiser, as it cannot be written as the gradient of a smooth (Lipschitz gradient) function.
A similar line of work learns regularizers explicitly [33, 20, 21, 27, 55, 24, 49, 39, 59, 28, 50], rather than implicitly through the denoiser. This approach benefits from direct interpretability of the regularizer and can inherit convergence guarantees from classical variational theory.
A complementary approach leverages the intrinsic statistical structure of the denoiser. A particularly notable example—and the subject of this work—is the MMSE denoiser under Gaussian noise. With the Tweedie/Stein identities [22, 60], the MMSE denoiser can be tied to the score function, endowing it with useful structural properties [45]. Further, [29] and [31] showed that the MMSE denoiser is $C^{\infty}$ and can be written as the proximal map of a regularizer (which is $C^{\infty}$) under various forward and noise models, and derived an explicit formula for this regularizer (24). This progress enabled the derivation of asymptotic convergence for both the ADMM [53] and PGD [65] variants of PnP using an MMSE denoiser. However, a gap in the theory remains: The explicit formula offers little insight into the behavior of the induced regularizer, and there are no nonasymptotic convergence rates tailored to MMSE denoisers. Table I highlights several closely related results on PnP-PGD, the assumptions they impose on the denoiser, and the convergence guarantees they establish.
I-D Contributions
We make the following contributions.
  • Characterization of the implicit regularizer of the MMSE denoiser. We show in Theorem III.2 that the MMSE denoiser is the proximal map of a regularizer that admits an explicit form as an upper Moreau envelope, i.e., $\phi_{\mathrm{MMSE}}(x)$ is, up to a constant,
    $\sigma^{2}\sup_{z}\mathopen{}\mathclose{{\left\{f_{Z}(z)-\tfrac{1}{2\sigma^{2}}\|z-x\|^{2}}}\right\},$ (8)
    on its domain, where $f_{Z}=-\log p_{Z}$ is the negative log marginal distribution from Eq. 2. This is a novel explicit form of the implicit regularizer in the MMSE denoiser.
  • Nonasymptotic convergence of MMSE-PGD via weak convexity. In Theorem III.2, we prove that $\phi_{\mathrm{MMSE}}$ is $1$-weakly convex on its domain. With this property we establish (to the best of our knowledge) the first nonasymptotic convergence guarantees for PGD-PnP with an MMSE denoiser in Theorem III.3 and Section III-B. In particular, we show that, under an $L$-smooth data-fidelity term $f$ with $L<1$, both $\|\nabla F(x_{k})\|$ and $\|x_{k+1}-x_{k}\|$ decrease sublinearly in terms of the best iterate.
  • Experiments. In Section IV, we consider the following experimental setups: (i) A one-dimensional synthetic example validating the upper Moreau envelope form of the implicit regularizer in the MMSE denoiser by comparing a learned DNN denoiser to its explicit form, (ii) Gaussian deblurring on the MNIST dataset [40], and (iii) computed tomography on the MayoCT dataset [44]. In both imaging tasks we observe the predicted sublinear convergence.
II Preliminaries
In this section, we collect some definitions and results used in the remainder of the paper.
Definition II.1 (Lower Moreau Envelope).
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a function with $\gamma>0$. The lower Moreau envelope of $f$ is given by
$M_{\gamma}f(x)\coloneqq\inf_{y}\mathopen{}\mathclose{{\left\{f(y)+\tfrac{1}{2\gamma}\|y-x\|^{2}}}\right\}.$ (9)
Definition II.2 (Upper Moreau Envelope).
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a function with $\gamma>0$. The upper Moreau envelope of $f$ is given by
$M^{\gamma}f(x)\coloneqq\sup_{y}\mathopen{}\mathclose{{\left\{f(y)-\tfrac{1}{2\gamma}\|y-x\|^{2}}}\right\}.$ (10)
Note that the lower Moreau envelope is the usual Moreau envelope, but we make the distinction in this paper since we use both lower and upper Moreau envelopes. For the infimum defining the lower Moreau envelope to be achieved, $f$ must be lower-semicontinuous. An important property of the lower Moreau envelope $M_{\gamma}f$ is that it shares global minimizers with $f$ whenever $f$ is lower-semicontinuous, i.e. $\operatorname*{arg\,min}_{x}f(x)=\operatorname*{arg\,min}_{x}M_{\gamma}f(x).$ See [7] for further background.
Definition II.3 (Weak convexity).
Let $f:\mathbb{R}^{n}\to\mathbb{R}$, $\gamma>0$. $f$ is called $\gamma$-weakly convex if the mapping $x\mapsto f(x)+\tfrac{\gamma}{2}\|x\|^{2}$ is convex.
Crucial to our analysis is the following lemma, adapted from [7, Lemma 3].
Lemma II.4.
Let $f:\mathbb{R}^{n}\to(-\infty,\infty]$. Then, $M^{\gamma}M_{\gamma}f(x)\leq f(x)$ for all $x\in\mathbb{R}^{n}$ and $M^{\gamma}f$ is $\tfrac{1}{\gamma}$-weakly convex.
Proof.
By definition,
$M^{\gamma}M_{\gamma}f(x)$
$={}$ $\sup_{y}\inf_{z}\mathopen{}\mathclose{{\left\{f(z)+\tfrac{1}{2\gamma}\|z-y\|^{2}-\tfrac{1}{2\gamma}\|y-x\|^{2}}}\right\}$
$\leq{}$ $\sup_{y}f(x)=f(x).$ (11)
Next,
$M^{\gamma}f(x)+\tfrac{1}{2\gamma}\|x\|^{2}$
$={}$ $\sup_{y}\mathopen{}\mathclose{{\left\{f(y)-\tfrac{1}{2\gamma}\|x-y\|^{2}+\tfrac{1}{2\gamma}\|x\|^{2}}}\right\}$
$={}$ $\sup_{y}\mathopen{}\mathclose{{\left\{f(y)-\tfrac{1}{2\gamma}\|y\|^{2}+\tfrac{1}{\gamma}\langle x,y\rangle}}\right\},$ (12)
which is a supremum of affine functions, and hence convex. Therefore, $M^{\gamma}f$ is $\tfrac{1}{\gamma}$-weakly convex. ∎
We now turn to the proximal operator, a central tool in optimization. We shall rely on the fact that MMSE denoisers themselves admit a proximal representation.
Definition II.5 (Proximal operator).
Let $\gamma>0$, and $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{\infty\}$ be proper and lower-semicontinuous. The proximal operator of $f$ with parameter $\gamma$ is the mapping
[…]
Note that $\operatorname{prox}_{\gamma f}$ is often set-valued. By coercivity of the quadratic term, $\operatorname{prox}_{\gamma f}(x)$ is non-empty for every $x\in\mathbb{R}^{n}$. For a proper, lower-semicontinuous function $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{\infty\}$, the infimum defining the lower Moreau envelope is attained at every minimizer of the proximal problem; that is,
[…]
for all $x\in\mathbb{R}^{n}$. When $f$ is a convex, lower-semicontinuous, and proper function, it is well known that $\operatorname{prox}_{\gamma f}$ and $M_{\gamma}f$ are also related by the following formula due to Moreau [47]:
$\nabla M_{\gamma}f(x)=\tfrac{1}{\gamma}(x-\operatorname{prox}_{\gamma f}(x)).$ (15)
Although the proximal operator of a nonconvex function is generally set-valued, when $\operatorname{prox}_{\gamma f}$ is an MMSE denoiser, it is always single-valued and continuous, which is the setting of our main results. This motivates the following extension of the Moreau gradient identity.
Theorem II.6 (Extension of the Moreau Gradient Identity).
Let $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{\infty\}$ be proper and lower-semicontinuous. If $\operatorname{prox}_{\gamma f}(x)$ is single-valued and continuous for all $x$, then the gradient of the lower Moreau envelope exists and takes the form
$\nabla M_{\gamma}f(x)=\tfrac{1}{\gamma}(x-\operatorname{prox}_{\gamma f}(x)),\quad x\in\mathbb{R}^{n}.$ (16)
Proof.
Let $h(x,z)\coloneqq f(z)+\tfrac{1}{2\gamma}\|z-x\|^{2}$. We have for all $z\in\mathbb{R}^{n}$, $t\in\mathbb{R}$,
[…]
From the definition, $M_{\gamma}f(x)=\min_{z}h(x,z)$. Fix some $x_{0}\in\mathbb{R}^{n}$ where there is a unique minimizer $\{z_{0}\}=\operatorname{prox}_{\gamma f}(x_{0})$ by assumption. Take some $u\in\mathbb{R}^{n}$. There is some $\{z_{t}\}=\operatorname{prox}_{\gamma f}(x_{0}+tu)$ so that
$M_{\gamma}f(x_{0}+tu)$
$={}$ $h(x_{0}+tu,z_{t})$
$={}$ $h(x_{0},z_{t})+\tfrac{t}{\gamma}\langle x_{0}-z_{t},u\rangle+\tfrac{t^{2}}{2\gamma}\|u\|^{2}$
$\geq{}$ $M_{\gamma}f(x_{0})+\tfrac{t}{\gamma}\langle x_{0}-z_{t},u\rangle+\tfrac{t^{2}}{2\gamma}\|u\|^{2}.$ (18)
Subtracting and dividing through by $t$, we get
$\liminf_{t\to 0}\tfrac{M_{\gamma}f(x_{0}+tu)-M_{\gamma}f(x_{0})}{t}$
$\geq{}$ $\liminf_{t\to 0}\mathopen{}\mathclose{{\left(\tfrac{1}{\gamma}\langle x_{0}-z_{t},u\rangle+\tfrac{t}{2\gamma}\|u\|^{2}}}\right)$
$={}$ $\tfrac{1}{\gamma}\langle x_{0}-z_{0},u\rangle,$ (19)
where the last equality is by the assumption that $\operatorname{prox}_{\gamma f}$ is continuous. Next,
$M_{\gamma}f(x_{0}+tu)\leq h(x_{0}+tu,z_{0})$
$={}$ $M_{\gamma}f(x_{0})+\tfrac{t}{\gamma}\langle x_{0}-z_{0},u\rangle+\tfrac{t^{2}}{2\gamma}\|u\|^{2}.$ (20)
Again, subtracting and dividing through by $t$,
$\limsup_{t\to 0}\tfrac{M_{\gamma}f(x_{0}+tu)-M_{\gamma}f(x_{0})}{t}$
$\leq{}$ $\limsup_{t\to 0}\mathopen{}\mathclose{{\left(\tfrac{1}{\gamma}\langle x_{0}-z_{0},u\rangle+\tfrac{t}{2\gamma}\|u\|^{2}}}\right)$
$={}$ $\tfrac{1}{\gamma}\langle x_{0}-z_{0},u\rangle.$ (21)
Combining (II) and (II), we have that the directional derivative of $M_{\gamma}f$ at $x_{0}$ in direction $u$ satisfies $(M_{\gamma}f)^{\prime}(x_{0},u)=\tfrac{1}{\gamma}\langle x_{0}-z_{0},u\rangle$. Therefore, the gradient is
$\nabla M_{\gamma}f(x_{0})=\tfrac{1}{\gamma}(x_{0}-z_{0})=\tfrac{1}{\gamma}\mathopen{}\mathclose{{\left(x_{0}-\operatorname{prox}_{\gamma f}(x_{0})}}\right),$ (22)
which completes the proof. ∎
Many similar versions of this identity are well known in the literature, but are often formulated in terms of prox-regularity. We included a simple self-contained proof for completeness. For more detail, see [56, Theorem 13.37].
We next review key results on MMSE denoisers and state their connection to proximal mappings, beginning with Tweedie’s formula.
Theorem II.7 (Tweedie’s Formula).
Let $X\sim p_{X}$, $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ be independent random variables in $\mathbb{R}^{n}$, and let $Z=X+\epsilon$, $\psi_{\sigma}(z)\coloneqq\mathbb{E}[X|Z=z]$. Then,
[…]
where $p_{Z}=p_{X}*\mathcal{N}(0,\sigma^{2}I)$.
For a thorough introduction to Tweedie’s formula, we refer the reader to [22]. Note that Tweedie’s Formula makes no assumptions on the prior density $p_{X}$, which may even include discrete atoms. We recall the following result on MMSE denoisers due to Gribonval et al. [29, 31].
Theorem II.8 (Main result from [31] and [29, Corollary 1]).
Let $X\sim p_{X}$ and $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ be independent random variables in $\mathbb{R}^{n}$, and let $Z=X+\epsilon$. Assume $p_{X}$ is non-degenerate.222i.e., there is no pair $v\in\mathbb{R}^{n}$, $c\in\mathbb{R}$ such that $\langle X,v\rangle=c$, almost surely. The MMSE denoiser under (2) given by $\psi_{\sigma}(z)\coloneqq\mathbb{E}[X|Z=z]$ has the following properties:
  1. (i)
    It is one-to-one from $\mathbb{R}^{n}$ onto its image.
  2. (ii)
    It is $C^{\infty}$, and so is its inverse.
  3. (iii)
    It can be written as $\psi_{\sigma}(z)=\operatorname{prox}_{\phi_{\mathrm{MMSE}}}(z)$, where
$\phi_{\mathrm{MMSE}}(z)\coloneqq$
[…]
  1. where $f_{Z}=-\log p_{Z}$.
  2. (iv)
    $\phi_{\mathrm{MMSE}}$ is $C^{\infty}$ on $\psi_{\sigma}(\mathbb{R}^{n})$.
III Main Results
In this section, we present our theoretical contributions. We begin by introducing a new structural representation of the implicit regularizer underlying the MMSE denoiser. This characterization serves as the foundation for deriving novel nonasymptotic convergence guarantees.
III-A New Characterization of the MMSE Regularizer
It is known (Theorem II.8) that the MMSE denoiser is the proximal map of an infinitely differentiable penalty. Yet beyond this, the structure of the associated penalty $\phi_{\mathrm{MMSE}}$ remains unclear. The only known explicit expression of $\phi_{\mathrm{MMSE}}$ (24) is cumbersome and offers little analytical insight. In this section, we show that $\phi_{\mathrm{MMSE}}$ admits an upper Moreau envelope representation in terms of $f_{Z}$, yielding useful insights, namely $1$-weak convexity, which we will leverage to prove stronger convergence results for the PGD using an MMSE denoiser.
For the following, we will assume the setting of the denoising problem (2), letting $X\sim p_{X}$, $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ be independent random variables in $\mathbb{R}^{n}$, $Z=X+\epsilon$, so that $p_{Z}=p_{X}*\mathcal{N}(0,\sigma^{2}I)$. We will let $f_{Z}\coloneqq-\log p_{Z}$.
Lemma III.1.
The negative log marginal distribution from the denoising problem (2) can be written as
$f_{Z}=\tfrac{1}{\sigma^{2}}M_{1}\phi_{\mathrm{MMSE}}+C_{x_{0}},$ (25)
where $\phi_{\mathrm{MMSE}}$ is as in (24) and $C_{x_{0}}$ takes the form
[…]
for any choice of $x_{0}\in\mathbb{R}^{n}$.
Proof.
Pick $x\in\mathbb{R}^{n}$. By Theorem II.7 and Theorem II.8,
$x-\operatorname{prox}_{\phi_{\mathrm{MMSE}}}(x)$ $=-\sigma^{2}\nabla\log p_{Z}(x)$
$=\sigma^{2}\nabla f_{Z}(x).$ (27)
Combined with Theorem II.6 yields
$\sigma^{2}\nabla f_{Z}(x)=\nabla M_{1}\phi_{\mathrm{MMSE}}(x).$ (28)
Integrating gives the result. ∎
Theorem III.2.
The implicit regularizer in the MMSE denoiser can be written as
[…]
where $C_{x_{0}}$ is defined in (26).
Proof.
Pick some $x\in\mathbb{R}^{n}$. By Section II,
$\phi_{\mathrm{MMSE}}(x)\geq M^{1}M_{1}\phi_{\mathrm{MMSE}}(x),$ (30)
which, by the previous result, we can rewrite as
$\phi_{\mathrm{MMSE}}(x)$ $\geq M^{1}(\sigma^{2}f_{Z}-\sigma^{2}C_{x_{0}})(x)$
$=\sigma^{2}M^{\sigma^{2}}f_{Z}(x)-\sigma^{2}C_{x_{0}}.$ (31)
For the reverse inequality, take some $x\in\psi_{\sigma}(\mathbb{R}^{n})$. There exists some $z\in\mathbb{R}^{n}$ so that $x=\psi_{\sigma}(z)=\operatorname{prox}_{\phi_{\mathrm{MMSE}}}(z)$. The infimum defining $M_{1}\phi_{\mathrm{MMSE}}$ is achieved at $x$ (cf. (14)), so that
$\phi_{\mathrm{MMSE}}(x)+\tfrac{1}{2}\|x-z\|^{2}$ $=M_{1}\phi_{\mathrm{MMSE}}(z)$
$=\sigma^{2}f_{Z}(z)-\sigma^{2}C_{x_{0}},$ (32)
where the second line holds by Section III-A. Rearranging, we have
$\phi_{\mathrm{MMSE}}(x)$ $=\sigma^{2}\mathopen{}\mathclose{{\left(f_{Z}(z)-\tfrac{1}{2\sigma^{2}}\|x-z\|^{2}}}\right)-\sigma^{2}C_{x_{0}}$
$\leq\sigma^{2}M^{\sigma^{2}}f_{Z}(x)-\sigma^{2}C_{x_{0}},$ (33)
which completes the proof. ∎
We illustrate this result in Fig. 1 and Fig. 2 under a variety of mixture-of-Gaussian and mixture-of-Laplacian priors. An overview of the experimental details is provided in Section IV-A.
Figure 1: Top: Calculated implicit regularizer in the MMSE denoiser ($\phi_{\mathrm{MMSE}}$), negative log-marginal $f_{Z}$, and negative log-prior $f_{X}$ for three mixture-of-Gaussian priors under unit Gaussian noise. Bottom: The learned regularizer $\phi^{\mathrm{learned}}_{\mathrm{MMSE}}$, along with the same reference curves. Details in Section IV-A.
Figure 2: Top: Calculated implicit regularizer in the MMSE denoiser ($\phi_{\mathrm{MMSE}}$), negative log-marginal $f_{Z}$, and negative log-prior $f_{X}$ for three mixture-of-Laplacian priors under unit Gaussian noise. Bottom: The learned regularizer $\phi^{\mathrm{learned}}_{\mathrm{MMSE}}$, along with the same reference curves. Details in Section IV-A.
III-B Proximal Gradient Descent Convergence Results
By Theorem III.2, $\phi_{\mathrm{MMSE}}$ equals $\sigma^{2}M^{\sigma^{2}}f_{Z}$ on $\psi_{\sigma}(\mathbb{R}^{n})$. By Section II, $M^{\sigma^{2}}f_{Z}$ is $\tfrac{1}{\sigma^{2}}$-weakly convex. Hence, $\sigma^{2}M^{\sigma^{2}}f_{Z}$ is $1$-weakly convex. Therefore, $\phi_{\mathrm{MMSE}}$ is $1$-weakly convex on $\psi_{\sigma}(\mathbb{R}^{n})$. We leverage this to derive nonasymptotic convergence guarantees for PnP-PGD with MMSE denoisers.
Consider the iteration
[…]
where $\psi_{\sigma}$ is the MMSE denoiser under Gaussian noise of variance $\sigma^{2}$ and $f$ is an $L$-smooth data-fidelity term. The step size is fixed because it is implicit in the MMSE denoiser. The associated objective for these iterates is
[…]
where $\phi_{\mathrm{MMSE}}$ is the implicit regularizer in the MMSE estimator, discussed in Section III-A.
When the objective $F$ is smooth, the map $x\mapsto\nabla F(x)$ is continuous. Therefore, measuring convergence via the gradient norm $\|\nabla F(x_{k})\|$ is natural as small gradient norms imply closeness to a stationary point. For nonconvex problems, $\|\nabla F(x_{k})\|$ may oscillate even as it trends downward, so nonasymptotic convergence is typically stated in terms of the best iterate up until the current iteration [4]. Since $f$ is smooth and $\phi_{\mathrm{MMSE}}$ is $C^{\infty}$ on its domain (Theorem II.8), the objective $F$ is $C^{\infty}$ on its domain. We measure convergence via $\min_{i\leq K}\|\nabla F(x_{i})\|$ at each iteration $K$.
Theorem III.3 (Convergence of Eq. 34).
Let $f$ be proper, bounded from below and $L$-smooth with $L<1$. Then for $F$ in Eq. 35, the iterates of (34) satisfy the following:
  1. (i)
    The sequence $(F(x_{k}))$ is non-increasing and converges.
  2. (ii)
    The sequence has finite length, i.e., $\sum_{k=0}^{\infty}\|x_{k+1}-x_{k}\|^{2}<\infty$, and $\|x_{k+1}-x_{k}\|$ converges to 0 at rate
    $\min_{i\leq K}\|x_{i+1}-x_{i}\|=\mathcal{O}\mathopen{}\mathclose{{\left(\tfrac{1}{\sqrt{K}}}}\right).$ (36)
  3. (iii)
    All cluster points of $x_{k}$ are stationary points of $F$.
Proof.
We have that $\sigma^{2}M^{\sigma^{2}}f_{Z}-\sigma^{2}C_{x_{0}}$ is a globally-defined 11-weakly convex function that agrees with $\phi_{\mathrm{MMSE}}$ on $\psi_{\sigma}(\mathbb{R}^{n})$. Next, each iterate of (34) clearly satisfies $x_{k}\in\psi_{\sigma}(\mathbb{R}^{n})$ for $k\geq 1$. The application of [34, Theorem 1] with $\tau=1$, $M=1$, and $L_{f}<1$ yields the result. ∎
Corollary III.4.
Under the assumptions of Theorem III.3,
$\min_{i\leq K}\|\nabla F(x_{i})\|=\mathcal{O}\mathopen{}\mathclose{{\left(\tfrac{1}{\sqrt{K}}}}\right).$ (37)
Proof.
From Theorem II.8 and Theorem III.2,
$x_{k+1}=\operatorname*{arg\,min}_{z}G_{k}(z),$ (38)
where
$G_{k}(z)$ $\coloneqq\sigma^{2}(M^{\sigma^{2}}f_{Z}(z)-C_{x_{0}})$
$\quad+\tfrac{1}{2}\|z-(x_{k}-\nabla f(x_{k}))\|^{2}.$ (39)
By Section II, $\sigma^{2}M^{\sigma^{2}}f_{Z}$ is $1$-weakly convex. Hence, $G_{k}$ is convex and we can conclude that its minimizer satisfies first-order optimality, i.e., $\nabla G_{k}(x_{k+1})=0$. We can write this as $\nabla\phi_{\mathrm{MMSE}}(x_{k+1})+x_{k+1}-(x_{k}-\nabla f(x_{k}))=0,$ because $\phi_{\mathrm{MMSE}}(z)$ and $\sigma^{2}M^{\sigma^{2}}f_{Z}(z)-\sigma^{2}C_{x_{0}}$ agree when $z=x_{k+1}$. Therefore,
[…]
From this, we have that
$\|\nabla F(x_{k+1})\|$ $\leq\|x_{k}-x_{k+1}\|+\|\nabla f(x_{k+1})-\nabla f(x_{k})\|$
$\leq 2\|x_{k+1}-x_{k}\|.$ (41)
The result then follows from Theorem III.3 (ii). ∎
Remark III.5.
We note that it may seem that the requirement of $L<1$ in Theorem III.3 is restrictive, it turns out that any $L$-smooth data-fidelity term can be scaled appropriately to meet this condition. For example, in our computed tomography experiment (Section IV-D), the forward operator is a discretized Radon transform, which has an $L$-smooth $f$, where $L\approx 140$. Thus, we simply scale the data-fidelity term to obtain empirical results that verify our theoretical convergence rate.
Remark III.6.
The asymptotic convergence of PGD using an MMSE denoiser has previously been established for a range of step sizes $\gamma>0$. In particular, [65] showed that PnP-MMSE converges for any $\gamma\leq 1/L$. Though the MMSE denoiser can be written as $\operatorname{prox}_{\phi_{\mathrm{MMSE}}}$, we note that $\operatorname{prox}_{\gamma\phi{\mathrm{MMSE}}}$ cannot be evaluated for any $\gamma\neq 1$. Consequently, any change in the step size can be realized by a rescaling of the data-fidelity term. For this reason, we assume an implicit step size of $1$ and scale the Lipschitz constant of the data-fidelity term accordingly.
IV Experiments
In this section, we numerically verify the theoretical results presented in Section III. The code to reproduce our experiments is publicly available on GitHub. [^3] https://github.com/sparsity-group/pnpmmse All training and testing was done on an NVIDIA RTX PRO 6000 Blackwell Max-Q Workstation Edition GPU using the Ubuntu 24.04 Windows Subsystem for Linux.
Central to our experiments is the task of estimating the MMSE denoiser of a given distribution under Gaussian noise. Following the work of [24], we choose to model the MMSE denoiser as the gradient of a softplus deep neural network (DNN). We found that this architecture performed better than rectified linear unit (ReLU) alternatives. We also incorporate skip connections at each layer, as they have been shown to aid in image denoising tasks [43]. Finally, by Tweedie’s formula (Theorem II.7), the true MMSE denoiser can be written as the gradient of $\tfrac{1}{2}\|\cdot\|^{2}+\log p_{Z}$, where $p_{Z}$ is the marginal distribution under Gaussian noise. For this reason, our choice of modeling the MMSE denoiser as the gradient of a DNN is well-motivated.
TABLE II: DNN architectural details for inverse problems experiments (Section IV-C and Section IV-D). $H$ denotes the hidden dimension and $L$ denotes the number of layers.
Gaussian Deblurring Computed Tomography
Input size $28\times 28$ $512\times 512$
Dimensions $H\times 28\times 28,\ H\times 14\times 14,\ H\times 14\times 14,$ $H\times 512\times 512,\ H\times 256\times 256,\ H\times 256\times 256,\ H\times 128\times 128$
$H\times 7\times 7,\ H\cdot 7^{2}\times 1\times 1,\ 64\times 1\times 1$ $H\times 128\times 128,\ H\times 64\times 64,\ H\times 64\times 64,\ H\times 32\times 32,$
$H\times 16\times 16,\ 64\times 1\times 1,\ 64\times 1\times 1$
$L$ 66 $11$
$H$ $64$ $256$
Direct connections ($W_{k}$) 4 Conv2d ($3\times 3$ kernel), 2 Linear 9 Conv2d ($3\times 3$), 1 Conv2d ($16\times 16$ kernel), 1 Linear
Skip connections ($A_{k}$) 3 Conv2d, 1 Linear 8 Conv2d ($3\times 3$), 1 Conv2d ($16\times 16$ kernel)
Figure 3: Best-iterate gradient norm and iterate difference over $100$ iterations of PnP-PGD with an MMSE denoiser for the MNIST Gaussian deblurring experiment in Section IV-C, plotted on a log-log scale. Both quantities decay sublinearly, consistent with a $\mathcal{O}(1/\sqrt{K})$ rate.
IV-A Numerical Illustration of Theorem III.2
Figure 4: PnP-PGD reconstruction results on three selections from the MNIST dataset under Gaussian blur (Section IV-C). From left to right: clean reference, blurred and noisy input, and reconstruction after 100 iterations. PSNR gains are roughly 13131414 dB are achieved across all three examples.
Figure 5: Best-iterate gradient norm and iterate difference over $1000$ iterations of PnP-PGD with an MMSE denoiser for the MayoCT Computed Tomography experiment in Section IV-D, plotted on a log-log scale. Both quantities decay sublinearly, consistent with a $\mathcal{O}(1/\sqrt{K})$ rate.
We first consider the denoising problem Eq. 2 with simple 1-dimensional mixture-of-Gaussian and mixture-of-Laplacian priors, denoted by $p_{X}$, with additive zero-mean, unit-variance Gaussian noise $\epsilon$. In each case, $p_{Z}$ denotes the marginal distribution of $z$. We now describe the DNN architecture used to learn the MMSE denoiser. Let
$h_{0}$ $\coloneqq\sigma(W_{0}x),\quad W_{0}\in\mathbb{R}^{H},$
$h_{k+1}$ $\coloneqq\sigma(W_{k+1}h_{k}+a_{k}x+b_{k}),\quad k=0,\ldots,L-1,$ (42)
and $g_{\theta}(x)\coloneqq w_{\mathrm{out}}^{\mathsf{T}}h_{3}+a_{\mathrm{out}}x+b_{\mathrm{out}}$, where $W_{k}\in\mathbb{R}^{H\times H}$ are the weights, $a_{k},b_{k}\in\mathbb{R}^{H}$ are the skip connections and biases, and $w_{\mathrm{out}},a_{\mathrm{out}},b_{\mathrm{out}}\in\mathbb{R}$ are the output weights and biases.
We use a softplus activation function
[…]
with $\beta=10$. There are $L=4$ hidden layers, with a hidden dimension of $H=50$. In particular, we model our MMSE denoiser as the gradient of the DNN, i.e., $\psi_{\theta}\coloneqq\nabla g_{\theta}$. To learn the denoiser, we optimize the objective
[…]
with a batch size $B=2000$, $x_{i}\sim p_{X}$, and $z_{i}=x_{i}+\epsilon$, $\epsilon\sim\mathcal{N}(0,1)$. We consider the cases where $p_{X}$ is either a mixture-of-Gaussian or mixture-of-Laplacian distribution. We trained the DNNs for $500$ epochs with a learning rate $10^{-3}$ followed by $500$ epochs with a learning rate of $10^{-4}$.
As discussed in Section III, the MMSE denoiser $\psi_{\sigma}$ can be written as the $\operatorname{prox}_{\phi_{\mathrm{MMSE}}}$, where $\phi_{\mathrm{MMSE}}$ is the implicit regularizer. To approximate this, we take advantage of [30, Eq. (8)] and write
[…]
where we compute $\psi_{\theta}^{-1}$ numerically via least-squares. When $\psi_{\theta}$ is the bona fide MMSE denoiser, $\phi^{\mathrm{learned}}_{\mathrm{MMSE}}$ is the bona fide implicit regularizer [30].
Figure 6: PnP-PGD reconstruction result on three axial slices from the MayoCT dataset. From left to right: clean reference, noisy sinogram, filtered back projection of the noisy sinogram, and PnP-PGD output using an MMSE denoiser with early stopping. PSNR and SSIM are reported for the noisy and denoised images. Reconstructions achieve PSNR gains of roughly 15–20 dB over the noisy input, with SSIM improvements around $0.65$.
Finally, we can also exactly calculate the true MMSE denoiser, $\psi_{\mathrm{MMSE}}$, via Tweedie’s formula (Theorem II.7). From Eq. 14 and Section III-A, we have that
[…]
which allows us to solve for $\phi_{\mathrm{MMSE}}$ explicitly.
In Figs. 1 and 2 we compare the learned vs. the calculated implicit regularizers for three mixture-of-Gaussian and three mixture-of-Laplacian priors, along with the reference curve $f_{X}\coloneq-\log p_{X}$. Since $\phi_{\mathrm{MMSE}}$ is only determined up to the additive constant $C_{x_{0}}$ in Section III-A, we shift the curves vertically for visual comparison.
IV-B DNN Architecture for Inverse Problems
In this subsection, we now detail the DNN architecture for the inverse problems experiments that appear in Section IV-C (Gaussian deblurring) and Section IV-D (computed tomography). In both cases, we define $h_{1}\coloneqq\sigma(W_{1}x)$, and let
$h_{k}\coloneqq\sigma(W_{k}h_{k-1}+A_{k-1}\mathcal{I}_{k}(x)+b_{k}),$ (47)
where $k=2,\ldots,(L-1)$ and $\mathcal{I}_{k}(\cdot)$ denotes bilinear interpolation of $x$ so that the dimension match for matrix-vector multiplication. We let the scalar output be $g_{\theta}(x)$, defined as
$w_{\mathrm{out}}^{\mathsf{T}}\sigma(W_{L}\operatorname{vec}(h_{L-1})+A_{L}\operatorname{vec}(\mathcal{I}_{L}(x))+b_{L}).$ (48)
Each layer is either a convolutional or fully-connected layer. There is a skip-connection at every layer except the first and last. The DNNs use a softplus activation Eq. 43 with $\beta=10$. We provide the exact architectural details in Table II. The DNN is trained with the mean-squared error loss Eq. 44 with $z_{i}=x_{i}+\epsilon$ for zero-mean Gaussian noise $\epsilon$ with variance $\sigma^{2}=0.04$.
IV-C Gaussian Deblurring
We next consider the Gaussian deblurring inverse problem Eq. 1, where $x$ is sampled from the MNIST data set of handwritten digits [40], uniformly at random. The forward model $A$ is a Gaussian blurring operator, computed by circular convolution with a $(3\times 3)$ Gaussian kernel with $\sigma^{2}=1$. The kernel is normalized to have unit mass. We normalize the blur operator to have operator norm <1<1 to satisfy the assumptions in Theorem III.3.
We use the DNN architecture described in Section IV-B, trained using the Adam optimizer with a cosine annealing learning rate schedule. The network is trained with batches of $500$ for 50 00050\,000 epochs with a final learning rate of $3\times 10^{-5}$. We let $y=Ax+\epsilon$, with $\epsilon$ zero-mean Gaussian noise with variance $\sigma^{2}=0.04$. We use the data-fidelity term $f(x)\coloneq\tfrac{1}{2}\|Ax-y\|^{2}$ to get iterates of the form Eq. 7. We plot $\min_{i\leq K}\|\nabla F(x_{i})\|$ for the first $100$ iterations in Fig. 3, which we computed via Eq. 40. We display the reconstruction results in Fig. 4. Observe that there is clear sublinear decay, consistent with our theoretical results in Theorem III.3 and Section III-B.
IV-D Computed Tomography
We now consider the computed tomography inverse problem. The forward operator $A$ is given by a parallel beam Radon transform discretized via the Operator Discretization Library (ODL) [1]. We scale the operation by $1/(1+\|A\|_{\mathrm{op}})$ in accordance with Theorem III.3. We use 200 projection angles, uniformly spaced over $0,\pi)$, with 400 detector bins, implemented using the Astra toolbox [63] with CUDA acceleration.
We use the DNN architecture described in Section IV-B, trained using the Adam optimizer with a cosine annealing learning rate schedule. The network is trained with batches of $64$ for 40 00040\,000 epochs with a final learning rate of $10^{-6}$. We use training and testing data from the MayoCT dataset [44] (resolution of $512\times 512$, pixel values ranging from 0 to 1). We let $y=Ax+\epsilon$, with $\epsilon$ zero-mean Gaussian noise with variance $\sigma^{2}=0.04$. We use the data-fidelity term $f(x)\coloneq\tfrac{1}{2}\|Ax-y\|^{2}$ to get iterates of the form Eq. 7. We plot $\min_{i\leq K}\|\nabla F(x_{i})\|$ for the first $1000$ iterations in Fig. 5, which we computed via Eq. 40. We display the reconstruction results for three slices of the MayoCT dataset in Fig. 6. Observe that there is clear sublinear decay, consistent with our theoretical result in Theorem III.3 and Section III-B.
V Conclusion
We presented a novel representation of the MMSE denoiser’s implicit regularizer in Theorem III.2 and used it to derive nonasymptotic convergence results for PnP-PGD using an MMSE denoiser in Theorem III.3 and Section III-B. Our analysis is strictly tied to the Gaussian-noise MMSE setting. Tweedie’s formula, for example, is not the same under other noise models [22], so we suspect this flavor of result is unique to the Gaussian setting. However, we expect analogous nonasymptotic guarantees to be attainable for other PnP schemes that use MMSE denoisers under Gaussian noise, such as PnP-ADMM or the proximal point method. Exploration of these extensions constitutes future work.
Acknowledgment
The authors would like to thank Stanislas Ducotterd and Guillaume Lauga for helpful feedback on this work.